If a vehicle goes from 0-160 mph in 1320 feet , what is the elapsed time it takes the vehicle to go that 1320 feet ,
Answer by Alan3354(69443)

If a vehicle goes from 0-160 mph in 1320 feet , what is the elapsed time it takes the vehicle to go that 1320 feet
--------------
Assuming uniform acceleration, its average speed over the distance is 80 mi/hr
---
d = r*t
t = d/r = (1/4)/80 = 1/320 of an hour